What is the present value of Rs 10,000 due in two years at 8% p.a,C.I according at the interest is paid (a) yearly or (b) half yearly


1
Expert's answer
2021-11-10T15:43:06-0500

a) "A=10,000""n=2"

"i=0.08"

Using the yearly means

"A=P(1+i)^n"

"10,000=P(1+0.08)^2"

"10,000=P(1.08)^2"

"10,000=1.1664P"

"P=Rs 8,573"

b) Half yearly

The formula is; "A=P(1+i\/2)^2n"

"10,000=P(1+0.08\/2)^4"

"10,00=P(1.04)^4"

"10,000=1.1699P"

"P=Rs8,548"